# Discovering Effective Policies for Land-Use Planning with Neuroevolution

### v5

@@ 3.2. Decision-making Problem
Bold: the best model in each region; *: 99% confidence the model outperforms LinReg; †: 99% confidence that RF outperforms NeuralNet, or vice-versa
4. Models
The system consists of the predictor, trained with supervised learning on the historical data, and the prescriptor, evolved against the predictor.
Prediction: Given the context and actions that were performed, the predictive model estimates the outcomes. In this case, since the cost outcome (i.e. the percentage of changed land) can be calculated directly, only the ELUC is predicted by the model. That is, given the land use of a specific location, and the land-use changes that were made during a specific year, the model predicts the long-term CO2 emissions directly caused by these land-use changes.
Any predictive model can be used in this task, including linear regression, random forest, or a neural network; they will each be evaluated below. As usual in machine learning, each model is fit to the existing historical data and evaluated with left-out data.
The optimal actions are not known, but the performance of each candidate strategy can be measured (using the predictive model), therefore the prescriptive model needs to be learned using search techniques. Standard reinforcement learning methods such as PPO and DQN are possible; the experiments in this paper use evolutionary optimization, i.e. conventional neuroevolution (Stanley et al., 2019). As in prior applications of ESP (Francon et al., 2020; Miikkulainen et al., 2021), the network has a fixed architecture of two fully connected layers; its weights are concatenated into a vector and evolved through crossover and mutation.
5. Experiments
This section describes the main results of the project so far, including different ways to create good predictors and the evolution of effective prescriptors based on them. The behavior of the discovered prescriptors is outlined, and an interactive demo of the entire land-use recommendation system is described.
5.1. Prediction
Second, a random forest model (RF) was trained on the same data but now including the full set of input features, in order to take advantage of RF’s ability to decide which features are most useful. The forest consisted of 100 decision trees with unrestricted maximum depth; at each split, a random subset of features was considered (where is the square root of the total number of features). To make the training time of the Global RF model feasible, it was trained with data from only 1982–2011. Preliminary experiments showed that constraining the training set in this manner did not decrease performance significantly. Both the linear regression and random forest models were trained using the scikit-learn library Pedregosa et al. (2011).
Third, a neural network (NeuralNet) was trained for the same task with the full set of features and data from 1851–2011 using the PyTorch library Paszke et al. (2019). The network was fully connected, composed of an input layer, a single hidden layer of size 4096, and an output layer. The output layer took as input the concatenation of the input layer and the hidden layer in order to take advantage of the observed linear relationships more easily (Cheng et al., 2016). It was trained for three epochs with batch size 2048 using the AdamW optimization algorithm Loshchilov and Hutter (2019). After each epoch, the learning rate was scaled by a factor of 0.1.
@@ 5.1. Prediction
LinReg, RF, and NeuralNet models were trained separately on each region and globally. While LinReg is deterministic and needed to be trained only once, RF and NeuralNet are stochastic in their initialization and training and were therefore trained 30 times to evaluate statistical significance. All models were tested on all regions with data from the most recent years, i.e. 2012–2021. As shown in Table 1, the LinReg models were fast to train, but they performed the worst (i.e., they have the largest mean absolute error in almost all cases), suggesting that the linear relationship is not enough to account for the data fully. RF performed significantly better than LinReg in all regions but EU, significantly worse than NeuralNet in EU and US, and outperformed NeuralNets in SA and globally. However, RF models do not generalize/extrapolate well to large changes in land use that occur in the dataset infrequently (or not at all for some land types; Figure 2), which turns out to be important when evolving prescriptors. Many prescriptor candidates suggest large land-use changes, and the predictor model needs to be able to evaluate them. In contrast, NeuralNets capture nonlinearities, which explains why they outperform LinReg in all regions. Most importantly, they extrapolate well, which suggests that they are a better choice for the surrogate model in prescriptor evolution. Additionally, LinReg and NeuralNet models have many fewer parameters than RF (taking 4KB and 363KB on disk respectively vs. 14GB for RF).
@@
In order to visualize the behavior of the predictor models, they were queried with a synthetic dataset created with every possible conversion from 100% land type A to 100% land type B (the types primf, primn, and urban were excluded from B as was explained in Section 3.2). Because the predictor takes latitude, longitude, cell area, and year as input, 24,204 (1%) latitude/longitude/cell-area/year combinations from the test dataset were sampled randomly and evaluated with each conversion. The results were averaged across the samples to get the overall result for the conversion. Figure 2 shows these results for the Global models. The first finding is that all models generally rank conversions to crops pasture forest in terms of reducing ELUC, and generally agree on the sign of ELUC. This consistency suggests that the results are reliable. Second, while LinReg and NeuralNet make similar predictions overall, their numerical values differ in many places, presumably due to the nonlinearities. Third, despite RF’s generally high accuracy in known scenarios, it is not able to generalize/extrapolate to large changes and instead predicts low-magnitude ELUC values for all imposed land-use transitions.
@@
Thus, since NeuralNet achieves similar or better performance than the RF in most cases, extrapolates well to extremes, and can learn nonlinear relationships in the data, the Global NeuralNet was used as the surrogate model to evolve the prescriptors.
5.2. Prescriptions
@@
The prescriptors were fully connected neural networks with two layers of weights, implemented with Keras/Tensorflow. The topology was fixed and only the weights were evolved. The input consisted of 12 float values, i.e. the eight land-use percentages and the latitude, longitude, and area of the geographical cell, as well as the year. The hidden layer contained 16 units with tanh activation.
@@
During evolution, prescriptor candidates were evaluated on a random subset of 1% of the Global dataset from years 1851–2011 with the Global NeuralNet predictor. Prescriptor output was a single vector of size five (the eight land-use percentages minus primf, primn, and urban, which were considered fixed), produced by the ReLU activation function. These five outputs were normalized to sum to the previous amount of land taken by these five types. The result is the prescriptor’s suggestion for the percentage of land use of each type. The five outputs were combined with the fixed land-use values for primf, primn, and urban to compute the difference from the initial land use. This difference was concatenated with the initial land use as well as latitude, longitude, cell area, and year, and passed to the Predictor model to get the predicted ELUC metric. This metric was aggregated across the evaluation data set to obtain the performance outcome of the suggestion. The cost outcome, i.e. the percentage of land-use change, was computed based on the difference between the five outputs and the corresponding inputs. All positive differences were summed for a given cell and divided by the total land-use percentage of the cell (excluding nonland).
@@
The aggregated ELUC and the land-use change outcomes were minimized with an evolutionary search on the neural network weights. Conventional neuroevolution with crossover and mutation operators on a vector of network weights was combined with NSGA-II (Deb et al., 2000) to account for multiobjectivity. To speed up evolution, two prescriptor models were trained with backpropagation and injected into the initial population: One to prescribe no land-use changes at all and the other to prescribe as much secondary forest (secdf) as possible (which was observed to reduce ELUC the most). The rest of the population was created randomly through an orthogonal initialization of weights in each layer with a mean of zero and a standard deviation of one (Saxe et al., 2014). Evolution was run for 100 generations with the following parameters:
nb_elites: 10
@@
mutation_factor: 0.2
@@
initialization_range: 1
mutation_probability: 0.2
remove_population_pct: 0.8
@@
Figure 3 demonstrates the progress of evolution towards increasingly better prescriptors, i.e. those that represent better implementations of each tradeoff of the ELUC and land-use change objectives. These solutions will be referred to as "Evolved Prescriptors" in the rest of the paper. They represent a wide variety of tradeoffs and a clear set of dominant solutions that constitute the final Pareto front. That set is returned to the decision-maker, who can then select the most preferred one to implement.
@@
To demonstrate the value of evolutionary search, the Evolved Prescriptors were compared against two heuristic baselines (Figure 4). Each of them converts as much land as possible to secdf, which has the lowest ELUC, up to a set threshold. The first is a simple baseline called "Even Heuristic", where an equal proportion of land up to the threshold is taken from all land-use types and converted to secdf. The second is an optimal linear baseline called "Perfect Heuristic", where as much land as possible is first taken from the land-use type that has the highest weight in the Global LinReg model, then from the land-use type that has the second highest weight, and so on until either the given land-use change threshold is reached or there is no land left to convert to secdf. One hundred such heuristic solutions were created with evenly distributed land-use change thresholds between 0% and 100% in order to generate heuristic Pareto fronts. All three methods were then evaluated on a random 1% sample of the Global test set from the years 2012–2021.
@@
As Figure 3 shows, the Evolved Prescriptors and the heuristics all converged in the high-change region of the Pareto front, where 100% forest prescriptions dominate. The heuristics performed slightly better in the low-change region, where the predictors may not be accurate enough to distinguish between small changes. However, in the most useful middle-change region, the Evolved Prescriptors outperformed both heuristics, thus demonstrating that they are likely exploiting the nonlinear relationships well. The performance for each prescription method can be quantified in terms of hypervolume, which is measured as the rectangular area above and to the right of each point (only counting overlapping area once), extending to ELUC = 0 and change = 1. The Even Heuristic is outperformed by the Perfect Heuristic which is outperformed by the Evolved Prescriptors, with hypervolumes of 19.68, 20.30, and 20.52, respectively.
5.3. Prescriptor Behaviors
@@
It is interesting to analyze how the Evolved Prescriptors were able to improve upon the two linear heuristics, which were constructed to be quite good to begin with. One representative from each method was chosen from the middle-change region and compared (Figure 5). The heuristic methods both prescribe 20% change and -12 and -14 ELUC on average, and the closest Evolved Prescriptor that dominates them averages 19% change and -16 ELUC. In Figure 5 these averages were expanded to show the results for actual samples in the test set (only 1% of those samples are plotted for readability). Note that the Evolved Prescriptor allocates more change to several samples than the Perfect Heuristic: This is because the heuristic’s change threshold is set to 0.35 (in order to get to 20% change on average) whereas the Evolved Prescriptor has no such restriction.
@@
The differences in change percentage and ELUC between the Perfect Heuristic and the Evolved Prescriptor are plotted for each sample in the test set in Figure 5. The heuristic dominates in the samples in the upper right quadrant (where both differences are above 0, i.e. the prescriptor’s change and ELUC are both larger than those of the heuristic) and the prescriptor in the bottom left quadrant. Surprisingly, the Evolved Prescriptor dominates in only 23 (0.09%) of the 24,204 test samples (and only one of them above the noise level, i.e. with the difference in ELUC greater than the MAE of the predictor). How can this be?
@@
The reason is that evolution discovered a way to take advantage of the large changes. In order to perform better on average, it does not need to outperform the Perfect Heuristic at the individual cell level; instead, it can identify and utilize cases where large changes make a big difference. Thus, even though the Evolved Prescriptor almost never beats the Perfect Heuristic in any individual cell, when averaged across all cells, it dominates the heuristic. In other words, evolution discovered the insight that it is most important to focus land-use changes in those regions where the largest changes are possible. Pick your battles right, and you’ll win overall.
@@
In order to see what the recommendations look like, the average policies of the three approaches are illustrated in Figure 6. The greatest amount of negative change made is to crop and positive is to secondary forest, i.e. all approaches recommend converting cropland to forest, which makes sense.
5.4. Interactive Evaluation

### v7

@@ 5.1 Prediction
Bold: the best model in each region; *: 99% confidence the model outperforms LinReg;
†: 99% confidence that RF outperforms NeuralNet, or vice-versa
@@
LinReg, RF, and NeuralNet models were trained separately on each region and globally. While LinReg is deterministic and needed to be trained only once, RF and NeuralNet are stochastic in their initialization and training and were therefore trained 30 times to evaluate statistical significance. All models were tested on all regions with data from the most recent years, i.e. 2012–2021, and evaluated on mean absolute error (MAE) of predicted ELUC. As shown in Table 1, the LinReg models were fast to train, but they performed the worst (i.e., they have the largest mean absolute error in almost all cases), suggesting that the linear relationship is not enough to account for the data fully. RF performed significantly better than LinReg in all regions but EU, significantly worse than NeuralNet in EU and US, and outperformed NeuralNets in SA and globally. However, RF models do not generalize/extrapolate well to large changes in land use that occur in the dataset infrequently (or not at all for some land types; Figure 2), which turns out to be important when evolving prescriptors. Many prescriptor candidates suggest large land-use changes, and the predictor model needs to be able to evaluate them. In contrast, NeuralNets capture nonlinearities, which explains why they outperform LinReg in all regions. Most importantly, they extrapolate well, which suggests that they are a better choice for the surrogate model in prescriptor evolution. Additionally, LinReg and NeuralNet models have many fewer parameters than RF (taking 4KB and 363KB on disk respectively vs. 14GB for RF).
@@
In order to visualize the behavior of the predictor models, they were queried with a synthetic dataset created with every possible conversion from 100% land type A to 100% land type B (the types primf, primn, and urban were excluded from B as was explained in Section 3.2). Because the RF and NeuralNet take latitude, longitude, cell area, and year as input, 24,204 (1%) latitude/longitude/cell-area/year combinations from the test dataset were sampled randomly and evaluated with each conversion. The results were averaged across the samples to get the overall result for the conversion. Figure 2 shows these results for the Global models. The first finding is that all models generally rank conversions to crops pasture forest in terms of reducing ELUC, and generally agree on the sign of ELUC. This ranking also generally aligns with those generated by the BLUE model. This consistency suggests that the results are reliable. Second, while LinReg and NeuralNet make similar predictions overall, their numerical values differ in many places, presumably due to the nonlinearities. Third, despite RF’s generally high accuracy in known scenarios, it is not able to generalize/extrapolate to large changes and instead predicts low-magnitude ELUC values for all imposed land-use transitions.
@@
Thus, since NeuralNet achieves similar or better performance than the RF, extrapolates well to extremes, and can learn nonlinear relationships in the data, the Global NeuralNet was used as the surrogate model to evolve the prescriptors.
5.2 Prescriptions
@@
The prescriptors were fully connected neural networks with two layers of weights, implemented with PyTorch. The topology was fixed and only the weights were evolved. The input consisted of 12 float values, i.e. the eight land-use percentages and the latitude, longitude, and area of the geographical cell, as well as the year. The hidden layer contained 16 units with tanh activation.
@@
During evolution, prescriptor candidates were evaluated on a fixed random subset of 1% of the Global dataset from years 1851–2011 with the Global NeuralNet predictor. Prescriptor output was a single vector of size five (the eight land-use percentages minus primf, primn, and urban, which were considered fixed), produced by the ReLU activation function. These five outputs were normalized to sum to the previous amount of land taken by these five types. The result is the prescriptor’s suggestion for the percentage of land use of each type. The five outputs were combined with the fixed land-use values for primf, primn, and urban to compute the difference from the initial land use. This difference was concatenated with the initial land use as well as latitude, longitude, cell area, and year, and passed to the Predictor model to get the predicted ELUC metric. The cost outcome, i.e. the percentage of land-use change, was computed based on the difference between the five outputs and the corresponding inputs. All positive differences were summed for a given cell and divided by the total land-use percentage of the cell (excluding nonland). These metrics were aggregated across the evaluation data set to obtain the performance outcome of each prescriptor candidate. Thus, a candidate was evaluated based on its average performance in creating a strategy for each land area in the test set.
@@
The aggregated ELUC and the land-use change outcomes were minimized with an evolutionary search on the neural network weights. Conventional neuroevolution with crossover and mutation operators on a vector of network weights was combined with NSGA-II [5] to make the search multiobjective. In order to seed evolution with good coverage of the search space, two prescriptor models were trained with backpropagation and injected into the initial population: One to prescribe no land-use changes at all and the other to prescribe as much secondary forest (secdf) as possible (which was observed to reduce ELUC the most). This process takes advantage of the Realizing Human Expertise through AI (RHEA) framework [21] for distilling human intuition into evolutionary search. Details and the impact of this seeding mechanism are discussed in Section 5.5. The rest of the population was created randomly through an orthogonal initialization of weights in each layer with a mean of zero and a standard deviation of one [32]. Evolution was run for 100 generations with the following parameters:
@@
mutation_factor: 0.1
mutation_probability: 0.2
@@
Figure 3 demonstrates the progress of evolution towards increasingly better prescriptors, i.e. those that represent better implementations of each tradeoff of the ELUC and land-use change objectives. These solutions are referred to as ”Evolved Prescriptors” in the rest of the paper. They represent a wide variety of tradeoffs and a clear set of dominant solutions that constitute the final Pareto front. That set is returned to the decision-maker, who can then select the most preferred one to implement.
@@
To demonstrate the value of evolutionary search, the Evolved Prescriptors were compared against two heuristic baselines (Figure 4). Each of them converts as much land as possible to secdf, which has the lowest ELUC, up to a set threshold. The first is a simple baseline called ”Even Heuristic”, where an equal proportion of land up to the threshold is taken from all land-use types and converted to secdf. The second is an optimal linear baseline called ”Perfect Heuristic”, where as much land as possible is first taken from the land-use type that has the highest weight in the Global LinReg model, then from the land-use type that has the second highest weight, and so on until either the given land-use change threshold is reached or there is no land left to convert to secdf. These heuristic solutions were created with evenly distributed land-use change thresholds between 0% and 100% in order to generate heuristic Pareto fronts. All three methods were then evaluated on a random 1% sample of the Global test set from the years 2012–2021.
@@
As Figure 4 shows, the Evolved Prescriptors and the heuristics all converged in the high-change region of the Pareto front, where 100% forest prescriptions dominate. The heuristics performed slightly better in the low-change region, where the predictors may not be accurate enough to distinguish between small changes. However, in the most useful middle-change region, the Evolved Prescriptors outperformed both heuristics, demonstrating that they are likely exploiting the nonlinear relationships well.
The performance of each prescription method can be quantified in terms of hypervolume, which measures how much of the solution space it dominates. It is calculated as the area above and to the right of the Pareto front up to maximum possible change and zero ELUC, divided by the area of the entire search space (which starts from zero change and lowest possible ELUC, which in turn is defined as the ELUC of the Perfect Heuristic at the maximum change). The Evolved Prescriptors outperform the Perfect Heuristic, which outperforms the Even Heuristic, with hypervolumes of 0.616, 0.566, and 0.512, respectively.
5.3 Prescriptor Behaviors
@@
It is interesting to analyze how the Evolved Prescriptors were able to improve upon the linear heuristics, which were constructed to be quite good to begin with. One representative from each method was chosen from the middle-change region and compared (Figure 5). The heuristic methods both prescribe 19.8% change and -12 and -14 ELUC on average, and the closest Evolved Prescriptor that dominates them averages 19.7% change and -18 ELUC. In Figure 5 these averages were expanded to show the results for actual samples in the test set (only 10% of those samples are plotted for readability). The samples for the Even Heuristic are hidden under those of the Perfect Heuristic: When these heuristics cannot reach their allocated thresholds, they change everything to secondary forest, resulting in overlapping points; their points similarly overlap at the vertical line indicating the change threshold. Note that the Evolved Prescriptor allocates more change to several samples than the Heuristic: This is because the heuristic’s change threshold is set to 0.35 (in order to get to 20% change on average) whereas the Evolved Prescriptor has no such restriction.
@@
The differences in change percentage and ELUC between the Perfect Heuristic and a typical Evolved Prescriptor are plotted for each sample in the test set in Figure 5. The heuristic dominates in the samples in the upper right quadrant (where both differences are above 0, i.e. the prescriptor’s change and ELUC are both larger than those of the heuristic) and the prescriptor in the bottom left quadrant. Surprisingly, the Evolved Prescriptor dominates the Perfect Heuristic in only 1 of the 24,204 test samples. How can this be?
@@
The reason is that evolution discovered a way to take advantage of large changes. As shown in Figure 5, the trained prescriptor is not constrained by a limit on every change it makes: As long as its changes remain below the limit on average, it can make a few large changes. As a result, evolution discovered a strategy where it makes only a few large changes, and makes them in cells where they have the largest impact. Thus, even though the Evolved Prescriptor almost never beats the Perfect Heuristic in any individual cell, when averaged across all cells, it dominates the heuristic. In other words, pick your battles right, and you’ll win overall.
@@
In order to see what the recommendations look like, the average policies of the three approaches are illustrated in Figure 6. The greatest amount of negative change made is to crop and positive is to secondary forest, i.e. all approaches recommend converting cropland to forest, which makes sense. The greatest departure from the heuristics is that the Evolved Prescriptors recommend more crop change and almost no rangeland change.
Where the Evolved Prescriptors picked their battles compared to the Perfect Heuristic is shown in Figure 6. Certain areas such as the Amazon rainforest were left untouched by both, because the system does not allow primary forests (that capture much carbon) to be changed. The Evolved Prescriptors avoided change in dry regions, such as the US Southwest, Sahara, Middle East, Central Asia, and Northwestern Australia, where presumably only limited change is effective. The prescriptors recommended more change in tropical, temperate, and continental regions like the US Midwest, Gulf of Guinea, South and Southeast Asia, Continental Europe, and the Pampas region of South America. These locations align with the change profile in Figure 6: The areas where more crops can grow are changed more, whereas plains and desert regions corresponding to rangeland are barely touched.
5.4 Constraining Crop Loss
While minimizing ELUC and change allows a decision-maker to find solutions along an important trade-off, it does not limit how much cropland the prescriptor is allowed to change, which could lead to food shortages. A step towards a more practical application is to optimize under such a constraint. To this end, the same evolutionary setup was run with a third objective of minimizing absolute change in crops. This approach assumes that the current level of crops is sufficient, however, the constraint could easily be set to an arbitrary value. Consequently, a seed model, trained through backpropagation, was included in the initial population to change as much land as possible to forest without removing any cropland. Such a seed makes it more likely for evolution to discover solutions that do not prescribe as much crop change. Evolution parameters were the same as in Section 5.2. These newly evolved prescriptor models were compared to the previous heuristics as well as to a new ”No Crop” Heuristic, i.e. the Perfect Heuristic without modifying any cropland.
Three main findings resulted from this experiment. First, the No Crop heuristic performs quite poorly in the other two objectives, demonstrating that more sophisticated optimization is necessary to incorporate this constraint. Second, evolution was able to find solutions trading off all three objectives, allowing a decision-maker to fix the amount of crop to change before selecting their preferred solution. Third, evolution was able to discover solutions that outperformed the Perfect Heuristic in all three objectives. The trade-off can be seen in Figure 7. The Pareto front is again plotted as in Figure 4, but this time, each Evolved Prescriptor is colored by the difference in how much it changed cropland on average compared to the Perfect Heuristic at the same level of change. There is a clear color gradient from more to less crop change perpendicular to the Pareto front such that generally the candidates that prescribed more crop change did the best and those that prescribed less did worse in terms of ELUC and change. Thus, a decision-maker can first choose how much crop they want to change, and then select the best prescriptor at that level along the Pareto front.
The solutions that outperform the Perfect Heuristic in all three objectives are shown in Figure 7 outlined in black. Each point represents a prescriptor that achieves a better ELUC with less change than the Perfect Heuristic while also prescribing less crop change. Figure 7 shows the average prescription of one such point in comparison to the heuristics in order to determine how it did so. The Evolved Prescriptor changed slightly less crop but compensated by removing more pasture. This strategy allowed it to prescribe less crop change, less change overall, and achieve a lower ELUC than the perfect heuristic, showing the power of evolution in the case of three objectives. Ultimately, the results of this experiment show the ability of the proof-of-concept system to incorporate realistic constraints such as those on food production in order to make it practical in real-world decision-making.
5.5 Incorporating Human Expertise
When developing solutions to global societal problems with AI, incorporating expert human knowledge may improve performance while also adding accountability and explainability [21]. The first step was taken in Section 5.2, where two hand-designed solutions were included in the initial prescriptor population through RHEA.
In more detail, RHEA consists of (1) defining the problem so that experts can develop solutions to it, (2) gathering a set of solutions from the experts, (3) distilling these solutions into a population of evolvable models such as neural networks, and (4) discovering better models by evolving this population. The land-use optimization problem was already defined with context, action, and outcome variables and was therefore compatible with the framework. Next, the behaviors for prescribing no change and all change to secondary forest were identified as useful extremes. In order to distill these behaviors, two sets of labels for 10% of the training dataset were created in order to run backpropagation. One was constructed by not changing the land use at all and the other changed as much land as possible to secondary forest. Then the AdamW optimization algorithm was run to minimize mean squared error over 15 epochs to targets that prescribed either no change or full change for a sampling of land-use cases. The resulting two trained prescriptors were then injected into the initial population and evolution was run.
In order to determine how useful it is in general to incorporate human expertise in this manner, two further experiments were run. One was an ablation study where no RHEA was performed and the entire initial population was randomly orthogonally initialized. The second one was an extended seeding study where the Perfect Heuristic was queried with thresholds from 5% to 95% in increments of 5%, prescriptors were trained with backpropagation to imitate the results of each query, and the resulting 19 seed models were added to the initial population together with the original two extreme prescriptors. The original and the new experiments were each run 10 times. Pareto fronts for each generation were created by combining all 10 trials’ candidates and then forming the Pareto front of the combined population. These Pareto fronts were also used to compute the hypervolume at each generation.
The results, shown in Figure 8, demonstrate the power of incorporating human expertise. When the distilled seed models were removed, even with 10 trials, evolution was unable to discover the top half of the Pareto front, missing out on the most realistic decision-making use cases (i.e. those with small changes). In Figure 8, the hypervolume of the No Seeding experiment barely reaches that of the Perfect Heuristic. However, just including one seed on each end of the Pareto front, as in the original seeded experiment, allowed evolution to achieve better results than the Perfect Heuristic. Additional seeds in the full RHEA case did not significantly improve the final solutions. Thus, simply injecting two extreme seeds based on human expertise was sufficient for this task.
The results of RHEA in a single run of the original two-seed experiment in Section 5.2 are further illustrated in Figures 8 and 8 (the results with other trials were qualitatively similar). Figure 8 shows the proportion of the final Pareto front that has a given candidate from the first generation as an ancestor. The no-change and full-change seed behaviors are ancestors of 94% and 87% of the final Pareto front, thus contributing to almost every prescriptor. The magnitude of that contribution is measured in Figure 8 by constructing a tree of parents for each prescriptor in the final Pareto Front and counting how many times each first-generation candidate was included in it. On average a model in the Pareto front is made up of over 30% of the distilled no-change behavior, mutated and crossed-over throughout the generations of evolution. This result aligns with the fact that the No Seeding experiment was unable to find the top left corner of the Pareto front: The seeded experiment needed to use the no-change solution frequently to fill in this top half. Very little behavior is taken from other randomly initialized models, showing the importance of the seeded model. Distilling this particular behavior was crucial in order to achieve the final Pareto front. The ultimate performance increase and explainability show how incorporating just a small amount of human feedback can improve the automatic decision-making process. In the future, as more real-world factors are taken into account, human expertise is likely to become even more valuable, and can be further applied with RHEA.
5.6 Interactive Evaluation